# A convex formulation of covariate-adjusted Gaussian graphical models via natural parametrization

### v1

@@ Acknowledgements
$\inf_{\boldsymbol{\mathbf{v}}\in\mathcal{V}}\frac{\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\boldsymbol{\mathbf{v}}\rVert_{2}}{\sqrt{n}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}}\geq\sqrt{\frac{m_{0}}{2}-\frac{1}{c_{0}}}>0.$
where $\mathcal{V}=\{\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p(q+1)-1}\mid\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}\rVert_{1}\leq 4\tau_{j}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}\}$
@@
$\geq\frac{m_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}.$
@@
$\geq\frac{m_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}-\frac{2}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}\rVert_{1}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{m_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}-\frac{8\tau_{j}}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{m_{0}}{2}-\frac{1+8\tau_{j}}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{m_{0}}{2}-\frac{(1+8\tau_{j})(s_{\beta}+s_{\gamma})}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}=\frac{m_{0}}{2}-\frac{1}{c_{0}}>0.$
In the last step we used that $c_{0}>2/m_{0}$ in Assumption 5. We have shown that
$\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}\geq\left(\frac{m_{0}}{2}-\frac{1}{c_{0}}\right)\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}^{2}$ (S25)
@@
$\frac{m_{0}}{2}\left\lvert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\ell}\right\rvert\leq\left\lvert\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right)_{\ell}\right\rvert+\frac{3}{c_{0}(1+8\tau_{j})(s_{\gamma}+s_{\beta})}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}.$
@@
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}\leq\frac{2}{m_{0}}\left\lVert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}+\frac{6}{c_{0}m_{0}(1+8\tau_{j})(s_{\gamma}+s_{\beta})}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}.$ (S26)
@@
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}^{2}\leq\frac{3}{2}(\lambda+\lambda_{g})\cdot(1+4\tau_{j})\left(\frac{2c_{0}}{c_{0}m_{0}-2}\right)\sqrt{s_{\gamma}+s_{\beta}}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}$
@@
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}\leq 3(\lambda+\lambda_{g})(1+4\tau_{j})\left(\frac{c_{0}}{c_{0}m_{0}-2}\right)\sqrt{s_{\gamma}+s_{\beta}}.$
@@
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}\leq 3(\lambda+\lambda_{g})(1+4\tau_{j})^{2}\left(\frac{c_{0}}{c_{0}m_{0}-2}\right)(s_{\gamma}+s_{\beta})$
@@
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}$ $\leq\frac{3}{m_{0}}(\lambda+\lambda_{g})+\frac{18(\lambda+\lambda_{g})(1+4\tau_{j})^{2}}{c_{0}m_{0}(1+8\tau_{j})}\left(\frac{c_{0}}{c_{0}m_{0}-2}\right)$
$=\frac{3}{m_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}m_{0}-2)}\right)$

### v2

@@ Acknowledgements
$\inf_{\boldsymbol{\mathbf{v}}\in\mathcal{V}}\frac{\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\boldsymbol{\mathbf{v}}\rVert_{2}}{\sqrt{n}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}}\geq\sqrt{\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}}}>0,$
where $\mathcal{V}=\{\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p(q+1)-1}\mid\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}\rVert_{1}\leq 4\tau_{j}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}\}$ and $c_{0}>2/\phi^{\ast}_{0}$ in Assumption 5.
@@
$\geq\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}.$
@@
$\geq\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}-\frac{2}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}\rVert_{1}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}-\frac{8\tau_{j}}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{\phi^{\ast}_{0}}{2}-\frac{1+8\tau_{j}}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{\phi^{\ast}_{0}}{2}-\frac{(1+8\tau_{j})(s_{\beta}+s_{\gamma})}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}=\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}}>0.$
We have shown that
$\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}\geq\left(\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}}\right)\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}^{2}$ (S25)
@@
$\frac{\phi^{\ast}_{0}}{2}\left\lvert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\ell}\right\rvert\leq\left\lvert\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right)_{\ell}\right\rvert+\frac{3}{c_{0}(1+8\tau_{j})(s_{\gamma}+s_{\beta})}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}.$
@@
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}\leq\frac{2}{\phi^{\ast}_{0}}\left\lVert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}+\frac{6}{c_{0}\phi^{\ast}_{0}(1+8\tau_{j})(s_{\gamma}+s_{\beta})}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}.$ (S26)
@@
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}^{2}\leq\frac{3}{2}(\lambda+\lambda_{g})\cdot(1+4\tau_{j})\left(\frac{2c_{0}}{c_{0}\phi^{\ast}_{0}-2}\right)\sqrt{s_{\gamma}+s_{\beta}}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}$
@@
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}\leq 3(\lambda+\lambda_{g})(1+4\tau_{j})\left(\frac{c_{0}}{c_{0}\phi^{\ast}_{0}-2}\right)\sqrt{s_{\gamma}+s_{\beta}}.$
@@
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}\leq 3(\lambda+\lambda_{g})(1+4\tau_{j})^{2}\left(\frac{c_{0}}{c_{0}\phi^{\ast}_{0}-2}\right)(s_{\gamma}+s_{\beta})$
@@
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}$ $\leq\frac{3}{\phi^{\ast}_{0}}(\lambda+\lambda_{g})+\frac{18(\lambda+\lambda_{g})(1+4\tau_{j})^{2}}{c_{0}\phi^{\ast}_{0}(1+8\tau_{j})}\left(\frac{c_{0}}{c_{0}\phi^{\ast}_{0}-2}\right)$
$=\frac{3}{\phi^{\ast}_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}\phi^{\ast}_{0}-2)}\right)$